find the area of a trapezium whose parallel sides are 7CM and 9cm long and whose altitude is 5cm.
Answer by mananth(16946)

Area of Trapezium = 1/2(sum of parallel sides) * height
A= 1/2 (7+9)*5
=1/2 * 16 *5
=40 cm^2